OpenStudy (republican31):

4 + x/3 - 2/3 = 7? Please help to solve, not just give the solution.

OpenStudy (anonymous):

start working on this bit by bit. you can take 4 off each side to begin with and it gives you x/3 -2/3 = 3

OpenStudy (anonymous):

remember - what you do to one side, you must do to the other. what would your next step be?

OpenStudy (republican31):

Okay. I was starting to go in the direction of converting the 4 into a fraction.

OpenStudy (republican31):

x-2/3 = 3?

OpenStudy (anonymous):

start with the simple addition and subtraction first. then deal with multiplication and division next

OpenStudy (republican31):

Next would I add 2 to both sides? I don't think that makes sense but I'm not sure. Or should I be dividing 3 by 3?

OpenStudy (anonymous):

close, but you'd do the 2/3. add 2/3 to each side.

OpenStudy (republican31):

x=3 2/3?

OpenStudy (anonymous):

\[4+\frac{ x }{ 3 }-\frac{ 2 }{ 3 }=7\]\[\frac{ x }{ 3 }-\frac{ 2 }{ 3 } = 3\]\[\frac{ x }{ 3 } = 3\frac{ 2 }{ 3 } = \frac{ 11 }{ 3 }\]\[x=11\]

OpenStudy (anonymous):

can you follow saseals working?

OpenStudy (republican31):

Up until the third line.

OpenStudy (anonymous):

3rd line you add 2/3 to both sides to cancel out the 2/3 on the x/3 side

OpenStudy (republican31):

Or nevermind, I see now.

OpenStudy (republican31):

Thank you both! But I can't give best response to both of you...

OpenStudy (anonymous):

you can give it to him

OpenStudy (republican31):

Alright! Thank you both very much! There might be a couple more questions coming from me tonight just so you know.